OpenStudy (anonymous):

solve: 6/5x+2/10x=28/10

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

12x + 2x = 28 x = 28/14 = 2

OpenStudy (anonymous):

(12+2)/10x =28/10 14/10x = 28/10 14 = 28x x =1/2

OpenStudy (anonymous):

multiply each side by 10 to get rid of ...

OpenStudy (anonymous):

thnk you everyone for yoour help!!!! You all ROCK!